# Large-eddy simulations of heat transfer from iced surfaces

### v1

@@ VI Heat transfer
In the rough region, the heat flux becomes highly heterogeneous. The highest values are typically observed on the windward sides of the roughness elements, where colder fluid impinges creating high-temperature gradients. The leeward sides exhibit slightly lower heat fluxes, likely due to flow separation and the formation of recirculating zones with reduced wall-normal temperature gradients. The tops of the roughness elements often show the lowest heat fluxes, even reaching negative values (heat is transferred from the fluid to the solid). As the roughness height increases (from case $s43$ to $s75$), the wall heat flux at the roughness peaks decreases, and a larger number of peaks exhibit negative heat fluxes. These results contrast with those from studies of rough isothermal walls, where the highest heat fluxes typically occur at the tops of the roughness elements due to their greater exposure to colder ambient air [45]. When thermal conduction within the solid is accounted for, however, this trend reverses. Although the tops of the elements are indeed exposed to colder fluid, they also become colder themselves due to limited thermal conduction from the base. To quantitatively characterize this effect, we define the spanwise average of a generic wall-variable $\phi_{w}$ (such as $\tau_{w}$, $T_{w}$ or $q_{w}$), conditioned at a specific vertical coordinate $y=y_{0}$, as follows:
[…]
@@
where $Z(x,y_{0})=\{z:f(x,z)=y_{0}\}$ represents the intersection between the surface $f(x,z)$ defining the rough wall and the horizontal plane at $y=y_{0}$. Figure 11 (left) shows the vertical profiles of the spanwise-averaged fluid temperature, $\langle T\rangle$, and the spanwise-averaged wall temperature conditioned by $y$, $\hat{T}_{w}$, at $x/L_{x}=0.6$ for geometry $s75$. The fluid temperature decreases gradually with wall-normal distance, whereas the wall temperature drops more rapidly. This indicates that, for low-conductivity materials, heat is transported more efficiently by convection in the fluid than by conduction within the solid. The reduction in the local temperature difference between the wall and the adjacent fluid with increasing wall-normal distance, shown in Figure 11 (right), closely correlates with the decrease in wall heat flux. This reduction in temperature difference underscores the critical role of internal conduction within roughness elements in controlling convective heat transfer over low-conductivity rough surfaces. Capturing these coupled mechanisms requires conjugate heat transfer simulations that resolve both solid and fluid thermal transport. The results presented in Figure 11 for $s75$ at $x/L_{x}=0.6$ are representative of the temperature profiles at other locations in the plate where roughness is present. Similar behaviours can be observed for $s43$ and $s55$. In these geometries, because roughness elements are shorter, the reduction in temperature difference is not as significant as for $s75$. This is reflected on the heat-flux maps where these two geometries have smaller regions with negative values of $q_{w}$.
@@
Figure 11: Left: Vertical profiles of the spanwise-averaged fluid temperature, $\langle T\rangle$, and the averaged wall temperature, $\hat{T}_{w}$, at $x/L_{x}=0.6$ for geometry $s75$. Right: Vertical distribution of the temperature difference between the fluid and the wall, along with the corresponding normalized wall heat flux.
VI.1 Stanton number distribution
@@
where $\Delta T=T_{w}\,-\,T_{\infty}$ is the difference between the surface temperature and the free-stream temperature. Measuring heat fluxes over rough surfaces presents significant challenges; therefore, to compute maps of $St$ (or $h$), the experiments assumed that $q_{w}=q_{\text{gen}}-q_{\text{HL}}-q_{\text{radiation}}$. We refer to this heat flux $q_{0}$, which has uniform values over each heating plate, and are equal to zero in between them. The Stanton number based on $q_{0}$ is denoted as $St_{0}$, which can be determined using $\Delta T$ from either the experiments ($St_{0,\textrm{exp}}$) or the simulations ($St_{0,\textrm{num}}$). However, this assumption for $q_{w}$ is not required in simulations, as the actual values of $q_{w}$ are directly accessible. Consequently, $St$ can be calculated using these local $q_{w}$ values without assuming uniform heat flux across the heating plates, and we designate the Stanton number calculated this way as $St_{\textrm{num}}$.
Figure 12: Spanwise-averaged $St$ profiles over a central band of thickness $\Delta z/L_{z}=1/2$ for the three geometries, obtained with the numerical model using a uniform heat flux distribution $q_{0}$ ($St_{0,\textrm{num}}$), as well as using the heat fluxes obtained from the simulation ($St_{\textrm{num}}$), superimposed with the $St$ obtained using the experimental temperatures and $q_{0}$ ($St_{0,\textrm{exp}}$). Discontinuities in black and orange lines correspond to the gaps between heating plates, where values of $q_{w}$ and $T_{w}$ were not reported in the experiments. Dashed line represents spanwise averaged surface elevation.
@@
VI.2 Comparison with isothermal walls
@@
Figure 13: Spanwise-averaged $St$ profiles for $s55$ geometry over a central band of thickness $\Delta z/L_{z}=1/2$, obtained with the numerical model using CHT ($St_{0,\textrm{num}}$ and $St_{\textrm{num}}$) and using isothermal walls.
VII Conclusions
@@
The examined cases exhibit large roughness-to-boundary-layer thickness ratios, ranging from $\kappa_{z}/\delta\approx 0.8$ where roughness is the highest ($X_{k,\max}$) to $\kappa_{z}/\delta\approx 0.2$ towards the end of the plate. These values indicate that the roughness effects penetrate the full boundary layer, disrupting outer-layer similarity and undermining the validity of empirical models based solely on surface statistics. As a result, such models tend to overpredict the equivalent sandgrain roughness ($k_{s}^{+}$) and wall-shear stress.
The simulations reveal that regions of high heat flux are concentrated on the windward faces of the roughness elements, while the leeward sides show moderately lower values. The crests of the roughness elements exhibit the lowest heat fluxes due to reduced temperature differences between the solid surface and the surrounding fluid. This distribution contrasts with that observed in simulations using isothermal walls, highlighting the effect of thermal conduction within the solid. For low-conductivity materials, such as ice or plastic, these results indicate that convective transport in the fluid is more efficient than conduction through the solid. The CHT simulations augment the experimental findings by enabling the calculation of $St$ accounting for spatial variations in heat flux that cannot be measured directly in experiments. These results underscore the importance of accounting for heat conduction when analyzing convective heat transfer over low-conductivity rough surfaces. Comparisons with isothermal wall simulations further reveal that neglecting solid conduction leads to an overprediction of $St$, particularly at the roughness crests.
@@ Appendix B Spanwise averaged velocity profiles
Figure 15 shows spanwise-averaged velocity deficit profiles for cases $s43$, $s55$ and $s75$. The velocity profiles for the three geometries do not collapse, indicating a loss of outer-layer similarity. Only in the downstream portion of the plates for cases $s43$ and $s55$, where the roughness height is relatively small, is outer-layer similarity partially recovered in the outer wake region. In contrast, case $s75$ exhibits significant variation between velocity profiles at all streamwise locations, showing that the influence of the roughness elements extends up to the edge of the boundary layer.

### v2

@@ VII Heat transfer
In the rough region, the heat flux becomes highly heterogeneous. The highest values are typically observed on the windward sides of the roughness elements, where colder fluid impinges creating high-temperature gradients. The leeward sides exhibit slightly lower heat fluxes. The tops of the roughness elements often show the lowest heat fluxes, even reaching negative values on the leeward side of the crest (heat is transferred from the fluid to the solid). As the roughness height increases (from case $s43$ to $s75$), the wall heat flux at the roughness peaks decreases, and a larger number of peaks exhibit negative heat fluxes (white areas in Figure 11). These results contrast with those from studies of rough isothermal walls, where the highest heat fluxes typically occur at the top of the roughness elements due to their greater exposure to colder ambient air [49]. When thermal conduction within the solid is accounted for, however, this trend reverses. Although the top of the elements are indeed exposed to colder fluid, they also become colder themselves due to limited thermal conduction from the base. This effect has been observed in previous studies [50, 51]; however, those studies did not report heat flux reversal at the roughness crests, likely because the roughness height considered was smaller. As our results show, increasing roughness height and accounting for low solid conduction can lead to stronger variation of heat fluxes, and even a local inversion of the heat flux direction.
Figure 12 presents the mean and instantaneous temperature fields in the solid and fluid regions of the domain around a roughness element, with mean velocity vectors superimposed. On the solid side, temperature decreases with height, with the windward face exhibiting lower temperatures than the leeward face due to enhanced cooling from the direct impact of the cooler fluid flow. Behind the roughness element, a recirculation zone forms, which entrains hot air from the base (which is in contact with the hottest parts of the surface) and transports it upward along the leeward side. This process leads to locally elevated temperatures on the leeward face. In some roughness elements, these local temperatures exceed the corresponding surface temperature, resulting in a reversal of the heat flux direction at the wall.
Figure 12: Mean (left) and instantaneous (right) temperature fields on the fluid and solid domain, superimposed with mean velocity vectors around a roughness element. Black line represents solid-fluid interface.
@@
To quantitatively characterize this effect, we compare spanwise-averaged profiles of wall temperature, $\langle T_{w}\rangle$, with spanwise-averaged profiles of fluid temperature, $\langle T\rangle$. It is important to note that, because we are considering rough surfaces, $\langle T_{w}\rangle=\langle T_{w}\rangle(x,y)$ varies with both the streamwise and wall-normal coordinates. Figure 13 (left) presents the vertical profiles of $\langle T\rangle$ and $\langle T_{w}\rangle$ at $x/L_{x}=0.6$ for geometry $s75$. The fluid temperature decreases gradually with wall-normal distance, whereas the wall temperature drops more rapidly. This indicates that, for low-conductivity materials, heat is transported more efficiently by convection in the fluid than by conduction within the solid. The reduction in the local temperature difference between the wall and the adjacent fluid with increasing wall-normal distance, shown in Figure 13 (right), closely correlates with the decrease in wall heat flux, showing that $\langle q_{w}\rangle\sim\langle T_{w}\rangle-\langle T\rangle$. This reduction in temperature difference underscores the critical role of internal conduction within roughness elements in controlling convective heat transfer over low-conductivity rough surfaces. Capturing these coupled mechanisms requires conjugate heat transfer simulations that resolve both solid and fluid thermal transport. The results presented in Figure 13 for $s75$ at $x/L_{x}=0.6$ are representative of the temperature profiles at other locations in the plate where roughness is present. Similar behaviours can be observed for $s43$ and $s55$. In these geometries, because roughness elements are shorter, the reduction in temperature difference is not as significant as for $s75$. This is reflected on the heat-flux maps where these two geometries have smaller regions with negative values of $q_{w}$.
@@
Figure 13: Left: Vertical profiles of the spanwise-averaged fluid temperature, $\langle T\rangle$, and the averaged wall temperature, $\langle T_{w}\rangle$, at $x/L_{x}=0.6$ for geometry $s75$. Right: Vertical distribution of the temperature difference between the fluid and the wall, along with the corresponding normalized wall heat flux. In this plot, $\langle\!\langle{\cdot}\rangle\!\rangle$ indicates the quantity averaged successively over the spanwise and wall-normal coordinates.
VII.1 Stanton number distribution
@@
where $\Delta T=T_{w}\,-\,T_{\infty}$ is the difference between the surface temperature and the free-stream temperature. Measuring heat fluxes over rough surfaces presents significant challenges; therefore, to compute maps of $St$ (or $h$), the experiments assumed that $q_{w}=q_{\text{gen}}-q_{\text{HL}}-q_{\text{radiation}}$, where $q_{\text{gen}}$, $q_{\text{HL}}$ and $q_{\text{radiation}}$ are the heat generated by the heating plate, the heat lost through the Plexiglass plate, and the heat lost due to radiation, respectively. We refer to this heat flux $q_{0}$, which has uniform values over each heating plate, and are equal to zero in between them. The Stanton number based on $q_{0}$ is denoted as $St_{0}$, which can be determined using $\Delta T$ from either the experiments ($St_{0,\textrm{exp}}$) or the simulations ($St_{0,\textrm{num}}$). However, this assumption for $q_{w}$ is not required in simulations, as the actual values of $q_{w}$ are directly accessible. Consequently, $St$ can be calculated using these local $q_{w}$ values without assuming uniform heat flux across the heating plates, and we designate the Stanton number calculated this way as $St_{\textrm{num}}$.
Figure 14: Spanwise-averaged StSt profiles over a central band of thickness $\Delta z/L_{z}=1/2$ for the three geometries, obtained with the numerical model using a uniform heat flux distribution $q_{0}$ ($St_{0,\textrm{num}}$), as well as using the heat fluxes obtained from the simulation ($St_{\textrm{num}}$), superimposed with the StSt obtained using the experimental temperatures and $q_{0}$ ($St_{0,\textrm{exp}}$). Discontinuities in black and orange lines correspond to the gaps between heating plates, where values of $q_{w}$ and $T_{w}$ were not reported in the experiments. Dashed line represents spanwise averaged surface elevation.
@@
VII.2 Comparison with isothermal walls
@@
Figure 15: Spanwise-averaged StSt profiles for $s55$ geometry over a central band of thickness $\Delta z/L_{z}=1/2$, obtained with the numerical model using CHT ($St_{0,\textrm{num}}$ and $St_{\textrm{num}}$) and using isothermal walls.
VIII Conclusions
@@
The examined cases exhibit large roughness-to-boundary-layer thickness ratios, ranging from $k_{z}/\delta_{99}\approx 0.8$ where roughness is the highest ($X_{k,\max}$) to $k_{z}/\delta_{99}\approx 0.2$ towards the end of the plate. These values indicate that the roughness effects penetrate the full boundary layer, disrupting outer-layer similarity and undermining the validity of empirical models based solely on surface statistics. As a result, such models tend to overpredict the equivalent sandgrain roughness ($k_{s}^{+}$) and wall-shear stress.
The simulations reveal that regions of high heat flux are concentrated on the windward faces of the roughness elements, while the leeward sides show moderately lower values. The crests of the roughness elements exhibit the lowest heat fluxes due to reduced temperature differences between the solid surface and the surrounding fluid. With increasing roughness height, the wall heat flux at the crest diminishes, even leading to heat flux reversal, where thermal energy is transferred from the fluid to the solid. This distribution contrasts with that observed in simulations using isothermal walls, highlighting the effect of thermal conduction within the solid. For low-conductivity materials, such as ice or plastic, these results indicate that convective transport in the fluid is more efficient than conduction through the solid. The CHT simulations augment the experimental findings by enabling the calculation of $St$ accounting for spatial variations in heat flux that cannot be measured directly in experiments. These results underscore the importance of accounting for heat conduction when analyzing convective heat transfer over low-conductivity rough surfaces. Comparisons with isothermal wall simulations further reveal that neglecting solid conduction leads to an overprediction of $St$, particularly at the roughness crests.
@@ Appendix B Spanwise averaged velocity profiles
Figure 17 presents spanwise-averaged velocity deficit profiles for cases $s43$, $s55$ and $s75$, compared to their smooth wall counterpart. The velocity profiles for all three roughness geometries fail to collapse onto the smooth wall profile, indicating a clear breakdown of outer-layer similarity. This deviation persists even in the downstream regions of the plate, where $k_{z}/\delta_{99}$ has diminished considerably. These results demonstrate that the influence of large roughness elements extends throughout the entire computational domain, affecting the flow field up to the boundary layer edge. This behavior is consistent with the observations reported by Hu et al. [52] for flow over a discrete roughness patch, where the effects of roughness extend a significant distance downstream of the end of the roughness patch.